# Non-destructive imaging of bulk electrical ‘hidden’ state switching in a 1-TaS2 cryo-memory device

### v1

@@ Methods
1-TaS2 single-crystals were synthesized using chemical vapor transport with iodine as the transport agent leading to consistent crystal quality [16].
Flakes were mechanically exfoliated from bulk crystals using GelPak and excess flakes removed using Scotch tape such that a large, single flake could be deposited onto an oxidized Si wafer (oxide thickness: ) with predefined Ti/Au alignment markers. The electrodes were written on the flake using electron beam lithography, followed by physical vapor deposition of Pd and Au. The investigated flake has lateral dimensions of about 50 and a thickness of . The in-plane (a,b) crystal orientation of the device was characterized at the Material Science beamline of Swiss Light Source synchrotron [61], and it was then glued onto a Cu sample holder using GE varnish.
Transport measurements and cooling procedure.
The device was mounted in a CryoVac 4He cryostat with Kapton windows for the incident and scattered X-rays. The resistance was measured in a two-probe configuration on the flake, and converted to four-probes on the sample holder such that only the resistance of the Pd/Au electrodes on the flake is measured in series with the flake resistance. The electrical setup consisted of a Keithley 6221 pulsed current source and a Keithley 2182 nano-voltmeter connected in delta mode. For electrical switching single, square-wave current pulses were applied using the current source. The sample temperature was monitored using a Cernox thermometer mounted on the sample holder about from the device. A specific cooling procedure was followed to ensure that the first-order phase transition to the CCDW state was crossed in a controlled fashion. The following ramp rates were used: 300 to : /, 250 to : /, 200 to : /, and 100 to : /.
X-ray diffraction and fluorescence. Spatially-resolved X-ray scattering was performed at the microXAS beamline of the Swiss Light Source [62], allowing for simultaneous in-situ resistance measurements, as well as XRD and XRF imaging. The incidence angle of the X-rays with respect to the device was fixed to , resulting in a beam spot size (v h) using Kirkpatrick-Baez focusing mirrors. The nominal X-ray flux was adjusted to about photons/s. Considering that the X-rays were traveling through about air before impinging on the sample, the effective incident flux on the sample was about and lower at 9.1 and , respectively. As the cryostat was mounted on xyz translation stages without rotational degree of freedom, to take XRD spatial maps we scanned through reciprocal space by changing the incoming X-rays energy between nd using the undulator and monochromator. The diffraction and fluorescence signals were recorded by Eiger X 4M and VIAMP-KC detectors, respectively. To locate the device the XRF signal at the Au () and Ta () -edges was used (see Supplementary Information).
@@
Data supporting the plots of this paper and other findings of this study are available from the corresponding authors upon request.
@@ Acknowledgements.
The authors acknowledge the Paul Scherrer Institute, Villigen, Switzerland for provision of synchrotron radiation at the microXAS and Material Science beamlines of the Swiss Light Source. We thank P. Sutar for synthesizing the 1-TaS2 crystals and acknowledge technical support from the PSI PICO operations team. Furthermore, we are grateful to B. Meyer, M. Birri and S. Stutz for technical support before and during the beamtimes. This research was funded by the Swiss National Science Foundation (SNSF) and the Slovenian Research And Innovation Agency (ARIS) as a part of the WEAVE framework Grant Number 213148 (ARIS project N1-0290). R.V., A.M. and D.M. thank ARIS for funding the research program P1-0400, and D.M. thanks ARIS for funding the research program J7-3146. B.L. and M.T. thank ARIS for funding the research program P2-0415. C.B., W.H., A.G.M. and G.A. acknowledge funding from the European Research Council under the European Union’s Horizon 2020 research and innovation program, within Grant Agreement 810451 (HERO). N.H. received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie Grant Agreement 884104 (PSI-FELLOW-III-3i).Author contributions
@@
Non-destructive imaging of bulk electrical ‘hidden’ state switching in a 1-TaS2 cryo-memory device
@@
Wenxiang Hu Henry G. Bell
Rok Venturini
Shih-Wen Huang
Aidan G. McConnell Faris Dizdarević
@@
Benjamin Lipovšek Marko Topič
Dimitrios Kazazis
Gabriel Aeppli
@@ I Fluorescence
X-ray fluorescence of Ta and Au was used to locate the device and set the range of the spatially-resolved scans. The signal was captured by VIAMP-KC detectors. Even small traces of elements such as Cu from the sample holder, Ti from the binary markers on the substrate, and Ni traces under the electrodes were detected (see Fig. S5). Fluorescence data were fitted using the PyMCA libraries [63].
@@ II.1 Experimental geometry constraint and reciprocal space reconstruction
The known parameters in this experiment are the energy of the X-rays, the orientation of the detector which was perpendicular to the X-ray beam and further detector specifications of the Eiger X 4M detector. In addition, we know the optimal pixel and energy of scattering reflections of three accessible lattice peaks (004), (013), (014) and four commensurate CCDW peaks on the detector and also their theoretical (h, k, l) parameters [48], leading to seven known quantities. The measured peak positions were determined by fitting a Lorentzian to the projected intensities. The CCDW peaks are elongated along the out-of-plane direction, therefore only the in-plane (h, k) dimensions were fitted. The unknown parameters in this setup were the relative position of the detector with respect to the sample (x, y, z) which was roughly measured and used as an initial guess, the sample orientation parameterized with angles () and the lattice constants of the flake (a, b, c). If we parametrize the Eiger X 4M orientation with Cartesian coordinates (x, y, z) where (x, y) point along the x, y- pixel direction and z perpendicular to the detector surface, the flake orientation can be expressed in this coordinate system using angles (). denotes the angle between the surface normal of the detector and the surface normal of the flake, is the azimuthal angle and the tilt angle. Since there were only seven measured parameters (three lattice and four CCDW reflections), we took the in-plane lattice parameters as known [65] and only the out-of-plane lattice parameter c as unknown.
@@
With this function, the reciprocal space coordinates of our measured peaks was compared to the theoretically known position [48, 28]. The cost function was minimized with respect to using a standard minimization algorithm from the SciPy optimization python library [66] which finds the experimental conditions of the unknown parameters that matches the experimental results. The resulting error table is shown in Tab. SII with small errors, meaning that there is good agreement between the measured peaks and the known structure of the material.
@@
The conversion from (, , E) to (h, k, l) leads to a curved surface in reciprocal space which was probed by the flat 2D detector in real space. Putting all the curved surfaces of the measured energies together, allowed us to reconstruct the 3D shape of the lattice, CCDW, HCDW and dimer peaks. Figure S6 shows this 3D reciprocal space representation measured in the fully-switched state (C). This is a volumetric plot using the intensities of the measured peaks, giving an overview over the peak shapes and position. Importantly, this is not a spatially-resolved 3D reciprocal image, it is the reconstruction coming from the sum of all pixels measured at the different energies, therefore we observe all peaks in the same plot. Next to the overview plot are extracts of reciprocal space around the lattice and CCDW peak in the unswitched state (A) and in the fully-switched state (C). We see that in the fully-switched state a HCDW peak appeared below the CCDW peak, as well as an additional lattice satellite (see Sect. V). For the spatially-resolved intensity maps of the device, we used such 3D reconstruction of every pixel. The peak positions were determined from the 3D reconstruction by summing along two reciprocal space directions to obtain a 1D intensity plot along the remaining reciprocal space direction and fitting a Lorentzian function to the peak.
@@ II.2 Gold ring correction
We observed five ring-like features in the raw diffraction patterns (see Fig. S7a). After the detector position with respect to the sample was determined, the angle of these rings, i.e. the radius , at a certain energy of the incoming X-rays was determined using Bragg’s law
where is the wavelength of the incoming X-rays, the spacing between the crystal planes, an integer for the diffraction order and the wavenumber. Comparison with literature showed that the radii of the rings matches Au with Miller indices (111), (200), (220), (311) and (222) [67] (see Fig. S7c). The electrodes are made from amorphous gold, therefore we observe these rings when spatially scanning over the device. Taking a region of interest (ROI) with radius ] (in m) corresponding to the thickness of the gold rings, allowed us to collect and sum up the intensity of all detector pixels within this ROI. We did this procedure for every pixel ( and ) on the sample which gave us spatially-resolved intensity maps outlining the electrodes of the device. We used these electrode images for aligning the spatial scans at the different X-ray energies.
After confirming that no peak of interest cut through the gold rings, we replaced them for the 3D reciprocal space reconstruction by an average background (see Fig. S7b). This was simply done for better visibility of the peak shapes in the reconstruction, but did not affect the reconstruction itself.
@@ II.3 Spatially-resolved maps
In the following, we describe how we obtain the spatially-resolved intensity maps, as well as the spatially-resolved projections in reciprocal space. For each state of the device, i.e. the room-temperature, as well as the unswitched (A), partially- (B) and fully-switched states (C) at the base temperature of 6 K, we measured spatial maps at different energies. As a next step, we performed a spatial jitter correction. That is, we spatially aligned the scans at different energies and for the different states. We did this by choosing a gold ring which shows clear outlines of the electrodes in the spatial maps. For each state, we first chose one reference spatial map measured at a certain X-ray energy. We then aligned all other spatial maps measured at different energies to this reference map. Due to the finite scan range, we confined a possible shift in the x and y spatial directions to and (in pixel) and found such that
@@
After having the scans at different energies for every state spatially aligned and a 3D reciprocal space reconstruction for each pixel performed, we defined 3D ROIs in reciprocal space around the peaks of interest. Since the CCDW and dimer peaks are elongated along l, as has been reported before [28], we took cylindrical ROIs around these peaks. For the lattice and HCDW peaks which are better defined along the out-of-plane direction, we took spherical ROIs. We computed the signal within these ROIs by summing up the intensities. For every spatial map at every energy, we set a spatial 1.5 (room temperature) and () grid according to the minimal step size of the scan. Then, we assigned intensity values to the grid by considering that if the closest scan to a respective grid pixel is 2 pixel, we set it to zero and otherwise we consider the data within the predefined ROI in reciprocal space. For the spatially-resolved intensity maps, such as in Fig. 2 of the main text, the intensity in the spatial pixels was therefore the sum of the intensities within the ROI. We note that with a minimal step size in the and direction of at room temperature and at the base temperature, we were in fact oversampling given the X-ray spot size of 1.5 (v h).
After identifying the region which switched to the HCDW state from the spatial maps, we could separate the device into different spatial regions to extract the spatially-resolved 2D reciprocal space projections (see Fig. 3 of the main text) and calculate the ratio of the HCDW and the CCDW peak intensity. The spatial region near the intergap space (light purple) corresponds to a intensity ratio and that region in the switching channel (dark purple) corresponds to a ratio of . Then, we summed up the 3D reciprocal space reconstructions of the pixels within these regions. For the projected maps, such as Fig. 3 of the main text, we set a grid in reciprocal space and interpolate. From this, we obtain the 2D projections by summing along one reciprocal space axis and dividing by the number of pixels in the spatial regions.
@@ III Room-temperature spatial maps
At the beginning of the experiments, the device was mounted at room temperature and spatial scans at different X-ray energies were measured to check the alignment, as well as to record the lattice and NCCDW signals. We took spatially-resolved scans at a few energies, corresponding to only selected cuts through reciprocal space. Hence, we only observed a few peaks with reduced statistics, compared to the low-temperature measurements. This resulted in a less accurate 3D reconstruction. Nevertheless, we took 3D ROIs around the (013) lattice peak and a respective NCCDW peak and plotted the corresponding spatial intensity map which both reproduce the general shape of the flake (see Fig. S8).
@@ IV Dimers and phase switching
As discussed in [28, 68], dimer peaks associated with a doubling of the real-space unit cell in the out-of-plane direction were observed in the CCDW state and vanish upon switching to the o-HCDW state [28]. We also observed dimer peaks in our measurements (see Fig. S6) and took 3D ROIs around those in reciprocal space. We recorded only a few energy scans cutting through the dimer peaks, meaning that the statistics was reduced compared to other peaks. Figure S9 shows the spatially-resolved intensity maps of the CCDW and dimer peak in the unswitched (A), the partially- (B) and the fully-switched state (C). We see that both peak intensities vanish in the bottom left part of the flake in the fully- switched state. Hence, we observe the same vanishing dimer peaks in the e-HCDW state in the intergap space as reported for the o-HCDW state, meaning for both there is also collapse of dimerization. Indirectly, observed as a local doubling of the out-of-plane dispersion, this has also been reported in a recent ARPES study [42]. As addressed in the main text, we observed that the intensity of the CCDW and dimer peaks in the unswitched state was the highest in the bottom left part of the flake that switches to the HCDW state. This supports the notion that the location of the switching channel may not only be dictated by the geometry of the device, but also the most homogeneous (dimerized) CDW order.

### v2

@@ Methods
1-TaS2 single-crystals were synthesized using chemical vapor transport with iodine as the transport agent leading to consistent crystal qualityĀ [19].
Flakes were mechanically exfoliated from bulk crystals using GelPak and excess flakes removed using Scotch tape such that a large, single flake could be deposited onto an oxidized Si wafer (oxide thickness:Ā ) with predefined Ti/AuĀ alignment markers. The electrodes were written on the flake using electron beam lithography, followed by physical vapor deposition of Ā Pd and Ā Au. The investigated flake has lateral dimensions of about 50 and a thickness of . The in-plane crystal orientation of the device was characterized at the Material Science beamline of Swiss Light Source synchrotron [65], and it was then glued onto a CuĀ sample holder using GEĀ varnish.
Transport measurements and sample cooling.
The device was mounted in a CryoVac 4HeĀ cryostat with Kapton windows for the incident and scattered X-rays. The resistance was measured in a two-probe configuration on the flake, and converted to four-probes on the sample holder such that only the resistance of the Pd/Au electrodes on the flake is measured in series with the flake. The electrical setup consisted of a Keithley 6221 pulsed current source and a Keithley 2182 nano-voltmeter connected in delta mode. For electrical switching single, square-wave pulses were applied using the current source. The sample temperature was monitored using a Cernox thermometer mounted on the sample holder about from the device. A specific cooling procedure ensured that the first-order phase transition to the CCDW state was crossed in a controlled fashion. The following ramp rates were used: 300 to :Ā /, 250 to : /, 200Ā toĀ : /, and 100 to : /.
X-ray diffraction and fluorescence. Spatially-resolved X-ray scattering was performed at the microXAS beamline of the Swiss Light SourceĀ [66], allowing for simultaneous in situ resistance measurements, as well as XRD and XRF imaging. The incidence angle of the X-rays with respect to the device was fixed to , resulting in a beam spot size (v h) using Kirkpatrick-Baez focusing mirrors. The nominal X-ray flux was adjusted to about Ā photons/s. Considering that the X-rays were passing through about air before impinging on the sample, the effective incident flux on the sample was about and lower at 9.1 and , respectively. As the cryostat was mounted on translation stages without rotational degree of freedom, to take XRD spatial maps we scanned through reciprocal space by changing the incoming X-rays energy between nd using the undulator and monochromator. The diffraction and fluorescence signals were recorded by Eiger X 4M and VIAMP-KC detectors, respectively. To locate the device the XRF signal at the Au () and Ta () -edges was used (Supplementary Information).
@@
Data supporting the figures and other findings of this study are available from the corresponding authors upon request.
@@ Acknowledgements.
The authors acknowledge the Paul Scherrer Institute, Villigen, Switzerland for provision of synchrotron radiation at the microXAS and Material Science beamlines of the Swiss Light Source. We thank P.Ā Sutar for synthesizing the 1-TaS2 crystals and acknowledge technical support from the PSIĀ PICO operations team. Furthermore, we are grateful to B.Ā Meyer, M.Ā Birri and S.Ā Stutz for technical support before and during the beamtimes, as well as fruitful discussions with N.Ā Taufertshƶfer. This research was funded by the Swiss National Science FoundationĀ (SNSF) and the Slovenian Research And Innovation AgencyĀ (ARIS) as a part of the WEAVE framework Grant NumberĀ 213148 (ARIS project N1-0290). R.V., A.M. and D.M. thank ARIS for funding the research program P1-0400, and D.M. thanks ARIS for funding the research program J7-3146. B.L. and M.T. thank ARIS for funding the research program P2-0415. C.B., W.H., A.G.M. and G.A. acknowledge funding from the European Research Council under the European Unionās Horizon 2020 research and innovation program, within Grant Agreement 810451 (HERO). N.H. received funding from the European Unionās Horizon 2020 research and innovation program under the Marie Sklodowska-Curie Grant Agreement 884104 (PSI-FELLOW-III-3i).Author contributions
@@
Imaging of electrically controlled van der Waals layer stacking in 1-TaS2
@@
Wenxiang Hu HenryĀ G.Ā Bell
RokĀ Venturini
Shih-WenĀ Huang
Aidan G. McConnell FarisĀ DizdareviÄ
@@
Benjamin LipovÅ”ek Marko TopiÄ
DimitriosĀ Kazazis
GabrielĀ Aeppli
@@ I Fluorescence
X-ray fluorescence of Ta and Au was used to locate the device and set the range of the spatially-resolved scans. The signal was captured by VIAMP-KC detectors. Even small traces of elements such as Cu from the sample holder, Ti from the binary markers on the substrate, and Ni traces under the electrodes were detected (Fig.Ā 5). Fluorescence data were fitted using the PyMCA librariesĀ [67].
@@ II.1 Experimental geometry constraint and reciprocal space reconstruction
The known parameters in this experiment are the energy of the X-rays, the orientation of the detector which was perpendicular to the X-ray beam and further detector specifications of the Eiger X 4M detector. In addition, we know the optimal pixel and energy of scattering reflections of three accessible lattice peaks (004), (013), (014) and four commensurate CCDW peaks on the detector and also their theoretical (h, k, l) parameters [51], leading to seven known quantities. The measured peak positions were determined by fitting a Lorentzian to the projected intensities. The CCDW peaks are elongated along the out-of-plane direction, therefore only the in-plane (h, k) dimensions were fitted. The unknown parameters in this setup were the relative position of the detector with respect to the sample (x,Ā y,Ā z) which was roughly measured and used as an initial guess, the sample orientation parameterized with angles () and the lattice constants of the flake (a, b, c). If we parametrize the Eiger X 4M orientation with Cartesian coordinates (x,Ā y,Ā z) where (x, y) point along the x, y pixel directions and z perpendicular to the detector surface, the flake orientation can be expressed in this coordinate system using angles (). denotes the angle between the surface normal of the detector and the surface normal of the flake, is the azimuthal angle and the tilt angle. Since there were only seven measured parameters (three lattice and four CCDW reflections), we took the in-plane lattice parameters as known [69] and only the out-of-plane lattice parameter c as unknown.
@@
With this function, the reciprocal space coordinates of our measured peaks was compared to the theoretically known position [51, 31]. The cost function was minimized with respect to using a standard minimization algorithm from the SciPy optimization python library [70] which finds the experimental conditions of the unknown parameters that matches the experimental results. The resulting error table is shown in Tab.Ā 2 with small errors, meaning that there is good agreement between the measured peaks and the known structure of the material.
@@
The conversion from (, , E) to (h, k, l) leads to a curved surface in reciprocal space which was probed by the flat 2D detector in real space. Putting all the curved surfaces of the measured energies together, allowed us to reconstruct the 3D shape of the lattice, CCDW, HCDW and dimer peaks. FigureĀ 6 shows this 3D reciprocal space representation measured in the fully-switched stateĀ C. This is a volumetric plot using the intensities of the measured peaks, giving an overview over the peak shapes and position. Importantly, this is not a spatially-resolved 3DĀ reciprocal image, it is the reconstruction coming from the sum of all pixels measured at the different energies, therefore we observe all peaks in the same plot. Next to the overview plot are extracts of reciprocal space around the lattice and CCDW peak in the unswitchedĀ A and in the fully-switched stateĀ C. We see that in the fully-switched state a HCDW peak appeared below the CCDW peak, as well as an additional lattice satellite (Sect.Ā V). For the spatially-resolved intensity maps of the device, we used such 3D reconstruction of every pixel. The peak positions were determined from the 3DĀ reconstruction by summing along two reciprocal space directions to obtain a 1D intensity plot along the remaining reciprocal space direction and fitting a Lorentzian function to the peak.
@@ II.2 Gold ring correction
We observed five ring-like features in the raw diffraction patterns (Fig.Ā 7a). After the detector position with respect to the sample was determined, the angle of these rings, i.e. the radius , at a certain energy of the incoming X-rays was determined using Braggās law
where is the wavelength of the incoming X-rays, Ā the spacing between the crystal planes, an integer for the diffraction order and the wavenumber. Comparison with literature showed that the radii of the rings matches Au with Miller indices (111), (200), (220), (311) and (222) [71] (Fig.Ā 7c). The electrodes are made from amorphous gold, therefore we observe these rings when spatially scanning over the device. Taking a region of interest (ROI) with radius ] (in m) corresponding to the thickness of the gold rings, allowed us to collect and sum up the intensity of all detector pixels within this ROI. We did this procedure for every pixel ( and ) on the sample which gave us spatially-resolved intensity maps outlining the electrodes of the device. We used these electrode images for aligning the spatial scans at the different X-ray energies.
After confirming that no peak of interest cut through the gold rings, we replaced them for the 3D reciprocal space reconstruction by an average background (Fig.Ā 7b). This was simply done for better visibility of the peak shapes in the reconstruction, but did not affect the reconstruction itself.
@@ II.3 Spatially-resolved maps
In the following, we describe how we obtain the spatially-resolved intensity maps, as well as the spatially-resolved projections in reciprocal space. For each state of the device, i.e. the room-temperature, as well as the unswitched A, partially- B and fully-switched statesĀ C at the base temperature of 6Ā K, we measured spatial maps at different energies. As a next step, we performed a spatial jitter correction. That is, we spatially aligned the scans at different energies and for the different states. We did this by choosing a gold ring which shows clear outlines of the electrodes in the spatial maps. For each state, we first chose one reference spatial mapĀ measured at a certain X-ray energy. We then aligned all other spatial maps measured at different energies to this reference map. Due to the finite scan range, we confined a possible shift in the x and y spatial directions to and (in pixel) and found such that
@@
After having the scans at different energies for every state spatially aligned and a 3D reciprocal space reconstruction for each pixel performed, we defined 3D ROIs in reciprocal space around the peaks of interest. Since the CCDW and dimer peaks are elongated along l, as has been reported before [31], we took cylindrical ROIs around these peaks. For the lattice and HCDW peaks which are better defined along the out-of-plane direction, we took spherical ROIs. We computed the signal within these ROIs by summing up the intensities. For every spatial map at every energy, we set a spatial 1.5 (room temperature) and () grid according to the minimal step size of the scan. Then, we assigned intensity values to the grid by considering that if the closest scan to a respective grid pixel is 2Ā pixel, we set it to zero and otherwise we consider the data within the predefined ROI in reciprocal space. For the spatially-resolved intensity maps, such as in Fig.Ā 2 of the main text, the intensity in the spatial pixels was therefore the sum of the intensities within the ROI. We note that with a minimal step size in the and direction of at room temperature and at the base temperature, we were in fact oversampling given the X-ray spot size of 1.5 (v h).
After identifying the region which switched to the HCDW state from the spatial maps, we could separate the device into different spatial regions to extract the spatially-resolved 2D reciprocal space projections (Fig.Ā 3 of the main text) and calculate the ratio of the HCDW and the CCDW peak intensity. The spatial region near the intergap space (light purple) corresponds to a intensity ratio and that region in the switching channel (dark purple) corresponds to a ratio of . Then, we summed up the 3D reciprocal space reconstructions of the pixels within these regions. For the projected maps, such as Fig.Ā 3 of the main text, we set a grid in reciprocal space and interpolate. From this, we obtain the 2D projections by summing along one reciprocal space axis and dividing by the number of pixels in the spatial regions.
@@ III Room-temperature spatial maps
At the beginning of the experiments, the device was mounted at room temperature and spatial scans at different X-ray energies were measured to check the alignment, as well as to record the lattice and NCCDW signals. We took spatially-resolved scans at a few energies, corresponding to only selected cuts through reciprocal space. Hence, we only observed a few peaks with reduced statistics, compared to the low-temperature measurements. This resulted in a less accurate 3D reconstruction. Nevertheless, we took 3D ROIs around the (013) lattice peak and a respective NCCDW peak and plotted the corresponding spatial intensity map which both reproduce the general shape of the flake (Fig.Ā 8).
@@ IV Dimers and phase switching
As discussed in [31, 72], dimer peaks associated with a doubling of the real-space unit cell in the out-of-plane direction were observed in the CCDW state and vanish upon switching to the o-HCDW state [31]. We also observed dimer peaks in our measurements (Fig.Ā 6) and took 3D ROIs around those in reciprocal space. We recorded only a few energy scans cutting through the dimer peaks, meaning that the statistics was reduced compared to other peaks. FigureĀ 9 shows the spatially-resolved intensity maps of the CCDW and dimer peak in the unswitched A, the partially- B and the fully-switched state C. We see that both peak intensities vanish in the bottom left part of the flake in the fully- switched state. Hence, we observe the same vanishing dimer peaks in the e-HCDW state in the intergap space as reported for the o-HCDW state, meaning for both there is also collapse of dimerization. Indirectly, observed as a local doubling of the out-of-plane dispersion, this has also been reported in a recent ARPES study [45]. As addressed in the main text, we observed that the intensity of the CCDW and dimer peaks in the unswitched state was the highest in the bottom left part of the flake that switches to the HCDW state. This supports the notion that the location of the switching channel may not only be dictated by the geometry of the device, but also the most homogeneous (dimerized) CDW order.